OpenStudy (isaiah_r30):

write three radical expressions that have 4√2s as their simplified form. What do the three expressions have in common? explain

OpenStudy (anthonyym):

If they simplify to the same thing, then that means they're all equal to one another. If a=z b=z c=z, then a = b = c

OpenStudy (isaiah_r30):

this stuff is confusing to me.

OpenStudy (anthonyym):

I think that's just the transitive property. Do you agree that they're equal? For example, if I have 2/4, 4/8, and 5/10, each simplifies to 1/2 so they are equal.

OpenStudy (isaiah_r30):

but dont they want me to have what simplifies to 4 sq 2

OpenStudy (anthonyym):

No, the question just asks "What do the three expressions have in common?"

OpenStudy (isaiah_r30):

It says write Three expression that have 4√2 as their simplified form

OpenStudy (isaiah_r30):

three radical*

OpenStudy (anthonyym):

Oh yeah misread my bad. So \[4\sqrt{2}=\sqrt{2*16}\] \[=\sqrt{32}\]

OpenStudy (anthonyym):

And you can make \[4\sqrt{2}\] into \[\frac{ 8\sqrt{2} }{ 2 }\]